# Continuous Audio Language Models

### v2

@@ 6 Conclusion
Appendix A Supplementary Experiments
A.1 Short-context transformer window
To study the influence of the context window $K$ of the short-context transformer, we performed an hyperparameter search with $K=5,10,20,40$ and trained our models for 500K training steps. We observe that finding an optimal value of $K$ is not critical even though it seems that 5 is not enough.
@@ A.1 Short-context transformer window
A.2 Scalability of CALM
In order to see if all the hyperparameters of our CALM method transfer well to more model parameters, we trained a consistency music CALM model with a larger 3B backbone (with a model dimension of 2048, 32 heads and 48 layers) as well as a discrete RQ-Transformer with 32-RVQ model with the same backbone size. For the CALM model, we use 4 consistency steps at inference. The results are as follow (we compare with the main model that has a 1.3B backbone):
@@ A.2 Scalability of CALM
Appendix B Human evaluation methods
@@ B.1 Speech continuation
B.2 Music continuation
@@ B.3 Bayesian Elo Score
which is a Bradley-Terry (Bradley & Terry, 1952) model. There are a few different methods to estimate the parameters of a Bradley-Terry model. We use the iterative one from (Caron & Doucet, 2010) where $S_{A}^{0}$ follows a Gamma prior with parameters $\alpha^{0},\beta^{0}$. By denoting $w_{A}$ the number of times where method $A$ won against any other methods and $n_{AB}$ the number of times where $A$ and $B$ are compared, $S_{A}^{t}$ is computed with the following update rule until convergence:
@@ Appendix C Hyperparameters
Table 9: Model and training hyperparameters
Music continuation Speech continuation
Backbone Transformer
Model dimension 1536 2560
MLP dimension 6336 10560
Number of heads 24 20
Number of layers 48 24
Learning rate $1\cdot 10^{-4}$ $5\cdot 10^{-5}$
Number of parameters 1.35B 2.2B
RQ-Transformer head (RVQ model)
Model dimension 1024 1024
MLP dimension 4096 4096
Number of heads 16 16
Number of layers 6 6
Number of parameters 701M 701M
Consistency sampler head (ours)
Number of layers 12 6
MLP dimension 3072 512
Gating SiLU SiLU
Number of parameters 601M 10M
Short context transformer (ours)
Model dimension 1536
MLP dimension 6336
Number of heads 24
Number of layers 4
Context 10
Number of parameters 113M
Audio embedding manipulations
Center and normalize
Noise before entering backbone
Training parameters
Head batch multiplier 8 8
Optimizer AdamW $\beta_{1}=0.9,\beta_{2}=0.95$ AdamW $\beta_{1}=0.9,\beta_{2}=0.95$
Batch size 48 144
Audio sample length 30s 300s
LR Schedule cosine cosine
Learning rate $1\cdot 10^{-4}$ $2\cdot 10^{-4}$
GPU used 16 H100 48 H100
Number of training steps 500k 150k
Initial checkpoint Helium-1 (Kyutai, 2025)
Inner monologue
Acoustic delay - 2 frames (160 ms)

### v3

@@ 6 Conclusion
7 Reproducibility statement
We will release the training code in order to retrain CALM. As well, we provide all the model hyperparameters in Sec. H.
Appendix A Lagrangian Self-Distillation
Lagrangian Self-Distillation (LSD) (Boffi et al., 2025) extends the consistency-model framework by introducing an additional time parameter $s$, enabling the model to learn mappings between arbitrary points along the probability flow trajectory rather than only from noisy inputs to clean data. An LSD model is a neural network $f_{\phi}(\mathbf{x},t,s)$ that predicts the state of the PF-ODE solution at time $s$ given its state $\mathbf{x}_{t}$ at time $t$: $\forall(s,t)\in[0,1]^{2},f_{\phi}(\mathbf{x}_{t},t,s)=\mathbf{x_{s}}$. In particular, with $t=1,s=0$ and by sampling $\mathbf{x}_{1}\sim\mathcal{N}(0,I)$ we can sample from the data distribution with $f_{\phi}(\mathbf{x}_{1},t=1,s=0)$.
In (Boffi et al., 2025), the authors define $\mathbf{F}_{\phi}(\mathbf{x},t,s)$ as
$f_{\phi}(\mathbf{x},t,s)=\mathbf{x}+(s-t)F_{\phi}(\mathbf{x},t,s).$ (4)
To train a LSD model from scratch we need to combine a flow matching loss as well as a LSD loss: $\mathcal{L}=\mathcal{L}_{FM}+\mathcal{L}_{LSD}$. With an adaptive weighting loss $w_{\psi}(t,s)$ and our notations they derive as:
$\mathcal{L}_{\text{FM}}(\phi,\psi)=\mathbb{E}_{\mathbf{x}_{0}\sim p_{\text{data}},\,\bm{\epsilon}\sim\mathcal{N}(0,\mathbf{I}),\,t\sim\mathcal{U}(0,1)}\left[e^{-w_{\psi}(t,t)}\left\|F_{\phi}(\mathbf{x}_{t},t,t)-\left(\alpha^{\prime}_{t}\mathbf{x}_{0}+\sigma^{\prime}_{t}\bm{\epsilon}\right)\right\|_{2}^{2}+w_{\psi}(t,t)\right].$ (5)
and
$\mathcal{L}_{\text{LSD}}(\phi,\psi)=\mathbb{E}_{\mathbf{x}_{0},\bm{\epsilon},t,s}\left[e^{-w_{\psi}(t,s)}\left\|\partial_{s}f_{\phi}(\mathbf{x}_{t},t,s)-F_{\phi^{-}}(f_{\phi}(\mathbf{x}_{t},t,s),s,s)\right\|_{2}^{2}+w_{\psi}(t,s)\right].$ (6)
where $\mathbf{x}_{0}\sim p_{\text{data}},\,\bm{\epsilon}\sim\mathcal{N}(0,\mathbf{I}),t\sim\mathcal{U}(0,1),s\sim\mathcal{U}(0,t)$.
In practice, we compute the flow matching loss on 75% of a batch and the LSD loss on the 25% left.
Appendix B Our Music VAE
EnCodec Copet et al. (2023) 4 RVQ 50 2.2 2.41 5.62
VQ-VAE (inspired from Mimi) 32 RVQ 25 8.8 3.63 9.61
VAE 32 dims 25 2.23 5.51
VAE 96 dims 25 3.65 9.76
VAE 128 dims 25 4.01 10.3
Our variational autoencoder (VAE) and codec architecture is adapted from the Mimi codec (Défossez et al., 2024b), originally designed for 24kHz speech at 12.5Hz. We trained it to compress 32kHz mono music with a 25Hz frame rate. We experiment with bottleneck sizes of 96 and 128 dimensions. For comparison, MusicGen’s EnCodec model (Copet et al., 2023) also operates at 32kHz but uses a 4-level RVQ at 50Hz. In Tab. 7, we report reconstruction metrics (audio ViSQOL Chinen et al. (2020) and SISNR), showing that a 32-dim VAE matches MusicGen’s codec, and that at least 96 dimensions are needed for our VAE to match the quality of the 32-level RVQ configuration.
Appendix C Effect of gaussian temperature sampling
We evaluate how our proposed gaussian temperature sampling method affects acoustic diversity for consistency models, in comparison with temperature sampling on its discrete counterpart. More precisely, we compute WavLM speaker embeddings (Chen et al., 2021b) over 100 unprompted speech generations with different values of temperature, both for an RQ-transformer and CALM. In both cases, the inner monologue text stream is generated with 0.8 temperature to ensure text diversity. We then average the pairwise cosine similarities of these embeddings, as a measure of diversity: the higher the average similarity, the lower speaker diversity in generated audio. The reference speaker similarity number is computed over 100 examples from the ground truth dataset. Fig. 2 shows, as expected, that speaker similarity tends to decrease with temperature, which means that temperature increases diversity with a similar trajectory to the discrete models.
Appendix D Data used for the text-to-speech model
The dataset used to train our TTS models is composed of AMI Carletta (2007), EARNINGS22 Del Rio et al. (2022), GIGASpeech Chen et al. (2021a), SPGISpeech O’Neill et al. (2021), TED-LIUM Hernandez et al. (2018), VoxPopuli Wang et al. (2021), LibriHeavy Beck et al. (2024), and Emilia Author and Contributors (2023). It results into 88k hours of audio.
Appendix E Supplementary Experiments
E.1 Head batch multiplier value
E.2 Short-context transformer window
To study the influence of the context window $K$ of the short-context transformer, we performed an hyperparameter search with $K=5,10,20,40$ and trained our models for 500K training steps (see Tab. 8). The model is the music continuation CALM and all evaluations are done with 4-steps of consistency. We observe that finding an optimal value of $K$ is not critical even though it seems that 5 is too small enough.
@@ E.2 Short-context transformer window
E.3 Comparison of TrigFlow and Consistency on inference speed and quality criteria
In Tab. 9 we compare the inference speed for both TrigFlow CALM and Consistency CALM on the task of music continuation on our test set. We see that below the 10 steps regime, only the Consistency model performs well. Moreover, around 10 steps is the limit for streaming applications where the real time factor (RTF) is smaller than 1. This justifies the need of consistency modeling (instead of diffusion or flow matching) for good quality streaming generation.
Table 9: Generation efficiency of TrigFlow CALM and Consistency CALM models. We compute the inference time to generate 30 seconds of audio, the corresponding Real Time Factor (RTF) as well as the FAD metric for different numbers of inference steps. For streaming (RTF<1), only consistency generates good quality audio.
# Steps Time (s) RTF FAD (TrigFlow) FAD (Consistency)
1 16.7 $0.56$ $0.83\pm 0.04$
4 20.4 $0.68$ $28.83\pm 0.20$ $0.71\pm 0.05$
10 27.7 $0.92$ $4.62\pm 0.07$ $0.73\pm 0.05$
25 46.1 $1.54$ $0.79\pm 0.04$ $0.96\pm 0.06$
50 76.8 $2.56$ $0.74\pm 0.05$ $1.46\pm 0.06$
100 136.4 $4.55$ $0.64\pm 0.04$ $2.05\pm 0.07$
E.4 Scalability of CALM
In order to see if all the hyperparameters of our CALM method transfer well to more model parameters, we trained a consistency music CALM model with a larger 3B backbone (with a model dimension of 2048, 32 heads and 48 layers) as well as a discrete RQ-Transformer with 32-RVQ model with the same backbone size. For the CALM model, we use 4 consistency steps at inference. Tab. 10 shows that the FAD are improving with a bigger backbone (3B vs 1.3B) in similar proportions for both discrete-based (RQ-Transformer) and continuous-based (CALM) models.
@@ E.4 Scalability of CALM
E.5 Text-to-Speech ablation study
Model Num. parameters WER CER SIM MUSHRA Speaker SIM
($\downarrow$) ($\downarrow$) ($\uparrow$) ($\uparrow$) (Human eval ELO ($\uparrow$)
CALM w/ LSD (NFE=1, CFG=1.5 w.r.t text) 313M 1.81 0.57 0.52 61.1 $\pm$ 2.3 1966 $\pm$ 23
CALM w/ LSD (NFE=1, no CFG) 313M 2.39 0.90 0.52 55.5 $\pm$ 2.2 1991 $\pm$ 26
CALM w/ LSD (NFE=1, CFG=1.25 w.r.t text and audio prefix) 313M 1.86 0.59 0.54 56.2 $\pm$ 2.2 1995 $\pm$ 27
CALM w/ Consistency (NFE=1, CFG=1.5 w.r.t text) 313M 1.93 0.62 0.40 46.8 $\pm$ 2.5 1900 $\pm$ 30
We can see in Tab. 10 that at this smaller scale ($\sim$300M), Lagrangian Self-Distillation models provide better audio quality than Consistency models. We also conclude that the latent CFG has a significant impact over the WER. As well, applying the latent CFG on the audio prefix improves the speaker similarity but at the cost of audio quality. All models were run with sampling from a Gaussian with 0.7 temperature (i.e., multiplying the standard Gaussian noise by $\sqrt{0.7}$).
E.6 Text-to-Music evaluation on MusicCaps
We report the results on the MusicCaps dataset Agostinelli et al. (2023) and report as well the results of MusicLM (Agostinelli et al., 2023), the original text-to-music MusicGen Medium (Copet et al., 2023), AudioLDM2 (Liu et al., 2024), Noise2Music (Huang et al., 2023), Jen-1 (Li et al., 2025) as well as MusicFlow (Prajwal et al., 2024) on the FAD, the KL Divergence (KLD) metric based on the pre-trained audio model PANN (Kong et al., 2019) as well as the CLAP cosine similarity that measures the matching between the generated music and text description. The results of Tab. 11 show that even though our goal is not to specifically build the best text-to-music model, simply applying CALM on our dataset with a CLAP conditioning leads to competitive results. The metrics from the models that we did not train are reported from the MusicFlow (Prajwal et al., 2024) paper.
Table 11: Text-to-music results on the MusicCaps dataset.
Model FAD ($\downarrow$) KLD ($\downarrow$) CLAP ($\uparrow$)
Reference 0.30
MusicLM (Agostinelli et al., 2023) 4.00
MusicGen Medium (Copet et al., 2023) 3.40 1.23 0.37
AudioLDM2 (Liu et al., 2024) 3.13 1.20 0.43
Noise2Music (Huang et al., 2023) 2.10
Jen-1 (Li et al., 2025) 2.00 1.29
MusicFlow (Unidirectional LM + FM) (Prajwal et al., 2024) 2.69 1.23 0.52
MusicGen Medium (retrained with CLAP) 2.70 1.37 0.39
RQ-transformer 32 RVQ 2.56 1.35 0.43
CALM - Consistency - 4 Steps 2.14 1.30 0.44
Appendix F Pocket TTS
Given a 313M parameters text-to-speech CALM model that has a 24 layers backbone transformer (the teacher), we do latent distillation to a 6 layers transformer backbone (the student) and a CFG coefficient of $\alpha=1.5$ applied to the teacher. We keep the same MLP sampling head. The final model, named Pocket TTS has a size of 90M parameters while the VAE has 20M parameters. In Tab.12, we put 100M parameters to include the VAE decoder in the parameter count.
We evaluate Pocket TTS on the Librispeech test-clean set following the same protocol as F5-TTS (Chen et al., 2025), with the difference that we cleaned the audio input using Adobe Enhance Speech [^1] https://podcast.adobe.com/en/enhance to obtain 24kHz high-quality audio. We evaluate all baselines with the enhanced samples [^2] https://huggingface.co/datasets/kyutai/librispeech_test_clean_enhanced.
We compare against three baselines: F5-TTS (Chen et al., 2025), DSM (Zeghidour et al., 2025), Chatterbox Turbo [^3] https://www.resemble.ai/chatterbox-turbo/ and Kokoro TTS [^4] https://huggingface.co/hexgrad/Kokoro-82M. We report Word Error Rate (WER) using Whisper-large-v3(Radford et al., 2022), as well as the results of a pairwise human evaluation for audio quality and speaker similarity. For audio quality, we ask raters “Which of the two audio clips has the best audio quality?”, and for speaker similarity, we ask “Which of the two audio clips sounds more similar to the reference audio clip in terms of voice characteristics?” and provide the voice prompt as a reference.
Table 12: Comparison of TTS models on Librispeech test-clean.
Model Param size WER Audio Quality Speaker Sim Faster than
(gen. only) ($\downarrow$) (ELO) ($\uparrow$) (ELO) ($\uparrow$) real-time CPU
F5-TTS (Chen et al., 2025) 336M 2.21 $1949\pm 27$ $1946\pm 26$ $\times$
DSM 750M 1.84 $1959\pm 25$ $2037\pm 21$ $\times$
Chatterbox Turbo 350M 3.24 $2055\pm 23$ $2012\pm 22$ $\times$
Kokoro 82M 1.93 No voice cloning $\checkmark$
Pocket TTS (ours) 100M 1.84 $2016\pm 25$ $1898\pm 26$ $\checkmark$
As seen in Tab. 12, Pocket TTS has the lowest Word Error Rate, a better Audio Quality than the ground truth F5-TTS and DSM, as well as an on-par Speaker Similarity with the ground truth while being a significantly smaller model than competitors, and being the only one that can run faster than real-time on CPU (we tested on Apple M3 and Intel core ultra 7 165H). We invite the reader to check the blog post where the one-line installation of the model is provided: kyutai.org/pocket-tts-technical-report.
Appendix G Human evaluation methods
@@ G.1 Speech continuation
G.2 Music continuation and CLAP-to-music
@@ G.3 Bayesian Elo Score
which is a Bradley-Terry (Bradley and Terry, 1952) model. There are a few different methods to estimate the parameters of a Bradley-Terry model. We use the iterative one from (Caron and Doucet, 2010) where $S_{A}^{0}$ follows a Gamma prior with parameters $\alpha^{0},\beta^{0}$. By denoting $w_{A}$ the number of times where method $A$ won against any other methods and $n_{AB}$ the number of times where $A$ and $B$ are compared, $S_{A}^{t}$ is computed with the following update rule until convergence:
@@ Appendix H Hyperparameters
Table 14: Model and training hyperparameters
Music Speech continuation Text-to-Speech
Backbone Transformer
Model dimension 1536 2560 1024
MLP dimension 6336 10560 4096
Number of heads 24 20 16
Number of layers 48 24 24
Learning rate $1\cdot 10^{-4}$ $5\cdot 10^{-5}$ 1e-4
Number of parameters 1.35B 2.2B 302M
RQ-Transformer head (RVQ model)
Model dimension 1024 1024
MLP dimension 4096 4096
Number of heads 16 16
Number of layers 6 6
Number of parameters 701M 701M
Consistency sampler head (ours)
Number of layers 12 6 6
MLP dimension 3072 512 512
Gating SiLU SiLU SiLU
Number of parameters 601M 10M 10M
Short context transformer (ours)
Model dimension 1536
MLP dimension 6336
Number of heads 24
Number of layers 4
Context 10
Number of parameters 113M
Audio embedding manipulations
Center and normalize
Noise before entering backbone
Training parameters
Head batch multiplier 8 8 8
Optimizer AdamW $\beta_{1}=0.9,\beta_{2}=0.95$ AdamW $\beta_{1}=0.9,\beta_{2}=0.95$ AdamW $\beta_{1}=0.9,\beta_{2}=0.95$
Batch size 48 144 128
Audio sample length 30s 300s 60s
LR Schedule cosine cosine cosine
Learning rate $1\cdot 10^{-4}$ $2\cdot 10^{-4}$ 1e-4
GPU used 16 H100 48 H100 8 H100
Number of training steps 500k 150k 400k
Initial checkpoint Helium-1 (Kyutai, 2025)
Inner monologue
Acoustic delay - 2 frames (160 ms) -